If 140 ml of 0.1 F AgNO3 is required to precipitate, quantitatively, silver chloride from 120 ml of KCl solution, what is the formality of the KCl solution the equation for this chemical reaction is Ag+ + Cl- =AgCl


1
Expert's answer
2021-03-02T01:01:33-0500

AgNO3 + KCl = AgCl + KNO3

n (AgNO3) = n (KCl)

n = CM x V

0.1 x 0.140 = CM(KCl) x 0.120

CM(KCl) = (0.1 x 0.140)/0.120 = 0.12 M